OpenStudy (anonymous):

Suppose $800 is invested into an account that pays 3.5% interest, compounded annually. What is the balance in the account after 5 years? A. $560.00 B. $950.15 C. $1142.86 D. $2800.00

OpenStudy (anonymous):

@KamiBug

OpenStudy (kamibug):

I think something is wrong with this question... Are you sure you wrote the right question and corresponding options?

OpenStudy (anonymous):

Suppose $800 is invested into an account that pays 3.5% interest, compounded annually. What is the balance in the account after 5 years? A. $560.00 B. $950.15 C. $1142.86 D. $2800.00

OpenStudy (kamibug):

Yeah, sorry I don't know.

OpenStudy (anonymous):

can u ask someone 4 me plz

OpenStudy (kamibug):

@mathmale @mathman @MathLanguage

OpenStudy (kamibug):

@SolomonZelman

OpenStudy (solomonzelman):

wait, so it's compounded ONCE EACH YEAR ?

OpenStudy (solomonzelman):

\[\huge\color{green}{ A=P(1+ \frac{r}{n})^{nt} } \] P = principal amount (the initial amount you borrow or deposit) r = annual rate of interest (as a decimal) t = number of years the amount is deposited or borrowed for. A = amount of money accumulated after n years, including interest. n = number of times the interest is compounded per year after I plug in numbers, I get \[\huge\color{blue}{ A=800(1+\frac{3.5 }{1} ) ^{ (1 \times 5)}}\]

OpenStudy (anonymous):

so is it a b c or d

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

plz help @wolf1728

OpenStudy (anonymous):

plz help @KyoshiKitty

OpenStudy (wolf1728):

$800 3.5 interest 5 years Total = Principal * (1+r)^years Total = 800 * (1.035)^5 Total = 800 * 1.1876863056 Total = 950.15

OpenStudy (anonymous):

i'll try @bubblebutt

OpenStudy (anonymous):

ii fan you and give u a medal if u can help me PPPPPPPPPLLLLLLLLLLLZZZZZZZZZZZZZZZ help meeeeeeeeeeeeeeeeeeeeeeeeeeee!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

sso b? @wolf1728

OpenStudy (wolf1728):

bubblebutt - that would be it

OpenStudy (anonymous):

u rock! :)

OpenStudy (wolf1728):

thanks - I'm good with compound interest and half life problems